# Large deviations of current for the symmetric simple exclusion process on a semi-infinite line and on an infinite line with a slow bond

### v2

@@ Abstract
The result (23) is for the specific initial density pair $(\rho_{a},\rho_{b})\equiv(1/2,0)$. For extending the result for other densities, we invoke a well-known rotational symmetry [31, 59, 58] of the least-Action (12). Essentially, the least-Action paths for two sets of parameters $(\lambda,\rho_{a},\rho_{b})$ and $(\lambda^{\prime},{\rho_{a}}^{\prime},{\rho_{b}}^{\prime})$ are related under a canonical transformation [31, 59, 58]. The symmetry results [38, 59, 58] in a dependence of the scgf on $(\lambda,\rho_{a},\rho_{b})$ through a single parameter
$\mu_{\text{si}}(\lambda,\rho_{a},\rho_{b})=R_{\text{si}}\big{(}\omega(\lambda,\rho_{a},\rho_{b})\big{)}$ (24)
where $\omega$ is defined in (2c). This dependence enables us [58] to deduce the function $R_{\text{si}}(\omega)$ from the result (23), leading to the expression (2b).
Slow bond: Recent interests [69, 70, 71, 72, 73, 35, 41, 54] in studying the effects of the coupling strength with reservoir can also be addressed in the semi-infinite line problem. The result (2b) is independent [54] of the coupling strength $\gamma$ (see Fig. 1) as long as it is larger than $\mathcal{O}(T^{-1/2})$. This is seen from the corresponding result for slow coupling $\gamma=\Gamma/\sqrt{T}$, where the boundary-fluctuations are significant, modifying the scgf (2b) to
@@
with $\sinh^{2}{u}=\omega(\lambda,\rho_{a},\rho_{b})$. In the $\Gamma\to\infty$ limit, (2b) is recovered from (25).
The expression (25) is obtained following an additivity argument [35], where contributions in (25) are separately from the single bond joining the reservoir and the system, and the system itself, optimised over the density at their common site [58]. This construction is very similar to the discussion in [35] for a related context, except for a crucial distinction that unlike the latter example, there is no quasi-stationarity for the semi-infinite problem. Nevertheless, the additivity conjecture gives the correct result (25), as verified in Fig. 4.
A similar additivity argument helps to solve the problem of current fluctuation across a single slow bond on the infinite lattice [74, 25, *2016_Franco_Phase, 75]. For the hopping rate $\gamma=\Gamma/\sqrt{T}$ across the slow bond, and unity for rest of the lattice, the generating function of current $\big{<}\mathrm{e}^{\lambda Q_{T}}\big{>}\sim\mathrm{e}^{\sqrt{T}\mu_{\text{inf}}^{\text{slow}}(\lambda,\rho_{a},\rho_{b})}$, for large $T$, with the scgf
@@
where $\sinh^{2}{u}=\omega(\lambda,\rho_{a},\rho_{b})$. In the $\Gamma\to\infty$ limit, the celebrated infinite line result [38, 39] is recovered [58] from (26).
Figure 4: The solid lines denote the scgf (25) for $\rho_{a(b)}=0.5$ and different values of $\Gamma$ (green for $\Gamma=1$, blue for $\Gamma=2$, and red for $\Gamma\to\infty$). The markers denote the corresponding numerical results obtained by the cloning algorithm with $T=500$ and $N_{c}=10^{4}$
@@
Other applications: Several other interesting conclusions can be drawn from our result. A known symmetry [31] of the MFT-Action extends our exact results to models [76, 77, 78] with a quadratic mobility $\sigma(\rho)=2A\rho(B-\rho)$, where $A$ and $B$ are arbitrary constants, culminating in $\mu_{\sigma}(\lambda,\rho_{a},\rho_{b})=(1/A)\,\mu_{\text{ssep}}\big{(}\omega(AB\lambda,\rho_{a}/B,\rho_{b}/B)\big{)}$.
@@
Our semi-infinite line result provides immediate solutions for two closely related problems. The first problem [24] concerns an infinite line SSEP with fast injection at a single site, a scenario which gained prominence in related contexts [79, 80, 81]. At long-times $t$, the net injection of particles $N_{t}$ on an empty lattice follows the asymptotics $\langle\mathrm{e}^{\lambda N_{t}}\rangle\sim\exp{[2\,\sqrt{t}\,\mu_{\text{si}}(\lambda,1,0)]}$ with (2b), which results from the statistical independence of the two halves due to fast injection. The second problem concerns the survival probability $P_{t}(k)$ of the $k^{\text{th}}$ particle (tracer) up to time $t$ in a fully packed SSEP on $\mathbb{Z}^{+}$ in the presence of an absorbing site at the origin. The asymptotics (7) and the relation $P_{t}(k)=P(Q_{t}>k)$ with the cumulative probability of integrated current result in a stretched exponential decay $P_{t}(k)\sim\exp{(-\sqrt{t/\tau_{s}})}$ for $k\lesssim 2\sqrt{t/\pi}$ at long-times $t$, with $\tau_{s}=[\phi(-k/\sqrt{t})]^{-2}$.
@@
Similar stretched exponential decays were noted in kinetically constrained models [82, 23, 83, 84], such as the energy-conserving spin-flip dynamics [23], where spin auto-correlation decays as $\mathrm{e}^{-\sqrt{t/\tau}}$, with $\tau$ remaining undetermined. For the spin-model, it was realized [84] that the domain-wall dynamics is equivalent to the infinite-line SSEP at a uniform density $\rho$. This correspondence relates  [^2] This work was initiated from a seminar by Mustansir Barma, who encouraged us to solve the problem. the spin-auto-correlation to the generating function of current in SSEP leading [58] to an exact result $\tau=[\mu(\mathrm{i}\pi,\rho,\rho)]^{-2}$, where $\mu(\lambda,\rho_{a},\rho_{b})$ is the scgf of the infinite line problem [38]. Similar correspondence holds in the presence of slow bond in the spin-flip dynamics, relating to (26).
@@
Conclusions: There remain several related open problems of immediate interest. Most prominent among them is the scgf for fixed (quenched) initial states at arbitrary densities. The only available non-trivial result [31] is for the infinite line SSEP at half-filling, obtained by establishing a relation with the fluctuating (annealed) initial state. A similar mapping for the semi-infinite line yields the quenched scgf
@@
Another pressing open question is the least-Action path at arbitrary densities for the semi-infinite line. In broader contexts, an extension of the problem in higher dimensions would be interesting, where only limited results are available [86]. From a practical point of view, the emergence of the infinite line SSEP in quantum circuits [8] is exciting, and a semi-infinite analog would be worth exploring. Along similar lines, extensions of our results for quantum analogues of the SSEP [87, 88] or for integrable models using generalized hydrodynamics [89, 90] would be timely.
@@ Acknowledgements.
We acknowledge the financial support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4002. TS thanks Bernard Derrida for insightful discussions on the problem. The importance of semi-infinite line SSEP and the additivity ansatz for slow bond originated from those discussions. TS also thanks Mustansir Barma for introducing him to the works on stretched exponential decays, particularly the open problem in [23].

### v3

@@ Abstract
The result (22) is for the specific initial density pair $(\rho_{a},\rho_{b})\equiv(1/2,0)$. For extending the result for other densities, we invoke a well-known rotational symmetry [2009_Derrida_Current2, 2010_Lecomte_Current, SM] of the least-Action (11). Essentially, the least-Action paths for two sets of parameters $(\lambda,\rho_{a},\rho_{b})$ and $(\lambda^{\prime},{\rho_{a}}^{\prime},{\rho_{b}}^{\prime})$ are related under a canonical transformation [2009_Derrida_Current2, 2010_Lecomte_Current, SM]. The symmetry results [2009_Derrida_Current, 2010_Lecomte_Current, SM] in a dependence (3a) of the scgf on $(\lambda,\rho_{a},\rho_{b})$ through a single parameter $\omega$ defined in (3d). This dependence enables us [SM] to deduce the function $R_{\text{si}}(\omega)$ from the result (22), leading to the expression (3b). Note that the $\omega$-dependence does not extend [2004_Derrida_Current, 2022_Mallick_Exact] for optimal profile $q_{\textrm{si}}(x,t)$, making it difficult to infer the profile at arbitrary densities.
Slow bond: Recent interests [2017_Baldasso_Exclusion, 2019_Franco_Non, 2020_Goncalves_Non, 2020_Erignoux_HydrodynamicsI, *2020_Erignoux_HydrodynamicsII, 2021_Derrida_Large, 2024_Saha_Large, 2023_Saha_Current] in studying the effects of the coupling strength with reservoir can also be addressed in the semi-infinite line problem. The result (3) is independent [2023_Saha_Current] of the coupling strength $\gamma$ (see Fig. 1) as long as it is larger than $\mathrm{O}(T^{-1/2})$. This is seen from the corresponding result for slow coupling $\gamma=\Gamma/\sqrt{T}$, where the boundary-fluctuations are significant, modifying the scgf (3a) to
@@
with $\sinh^{2}{u}=\omega(\lambda,\rho_{a},\rho_{b})$. In the $\Gamma\to\infty$ limit, (3a) is recovered from (23). The expression (23) is obtained following an additivity argument [2021_Derrida_Large], where contributions in (23) are separately from the single bond joining the reservoir and the system, and the system itself, optimised over the density at their common site [SM]. This construction is very similar to the discussion in [2021_Derrida_Large] for a related context, except for a crucial distinction that unlike the latter example, there is no quasi-stationarity for the semi-infinite problem. Despite this, the additivity conjecture gives the correct result (23).
A similar additivity argument helps to solve the problem of current fluctuation across a single slow bond at the origin on an infinite one-dimensional lattice [2013_Franco_Hydrodynamical, 2013_Franco_Phase, *2016_Franco_Phase, 2020_Erhard_Non], initially filled with uniform average density $\rho_{a(b)}$ in the negative (positive) half-line. For the hopping rate $\gamma=\Gamma/\sqrt{T}$ across the slow bond, and unity for rest of the lattice, the generating function of current $\big<\mathrm{e}^{\lambda Q_{T}}\big>\sim\mathrm{e}^{\sqrt{T}\mu_{\text{inf}}^{\text{slow}}(\lambda,\rho_{a},\rho_{b})}$, for large $T$, with the scgf [SM]
@@
where $\sinh^{2}{u}=\omega(\lambda,\rho_{a},\rho_{b})$. In the $\Gamma\to\infty$ limit, the celebrated infinite line result [2009_Derrida_Current, 2022_Mallick_Exact] is recovered [SM] from (24).
@@
Other applications: Several other interesting conclusions can be drawn from our result. A known symmetry [2009_Derrida_Current2] of the MFT-Action extends our exact results to models [1982_Kipnis_Heat, 2007_Giardinà_Duality, 2014_Vafayi_Weakly, 2022_Franceschini_Symmetric] with a quadratic mobility $\sigma(\rho)=2A\rho(B-\rho)$, where $A$ and $B$ are arbitrary constants, culminating in $\mu_{\sigma}(\lambda,\rho_{a},\rho_{b})=(1/A)\,\mu_{\text{ssep}}\big(\omega(AB\lambda,\rho_{a}/B,\rho_{b}/B)\big)$. Appropriate limits of $A$ and $B$ give the result for the Kipnis-Marchioro-Presutti (KMP) model [1982_Kipnis_Heat, 2007_Giardinà_Duality], the symmetric simple inclusion process (SSIP) [2014_Vafayi_Weakly, 2022_Franceschini_Symmetric], and the non-interacting particles [2023_Saha_Current].
@@
The semi-infinite line result provides solutions for two closely related problems. The first problem [2012_Krapivsky_Symmetric, 2014_Krapivsky_Lattice] concerns an infinite line SSEP with fast injection at a single site, a scenario which serves as the simplest description of hard-core particles spreading on an initially empty system [2019_Krapivsky_Free, *2020_Krapivsky_Free]. Such dynamics arise in diverse contexts, including the voter model [2003_Mobilia_vote], chemical catalysis [1992_Krapivsky_monomer, 1996_Krapivsky_catalytic], and wetting thin film spreading [Popescu_2012_wetting_films]. At long-times $t$, the net injection of particles $N_{t}$ on an empty lattice follows the asymptotics $\langle\mathrm{e}^{\lambda N_{t}}\rangle\sim\exp{[2\sqrt{t}\mu_{\text{si}}(\lambda,1,0)]}$ with (3), which results from the statistical independence of the two halves due to fast injection. The second problem concerns the survival probability $P_{t}(k)$ of the $k^{\text{th}}$ tagged particle (tracer) up to time $t$ in a fully packed SSEP on $\mathbb{Z}^{+}$ in the presence of an absorbing site at the origin. The problem has applications to first passage [2013_Majumdar_first_passage, 2014_Benichou_first_passage] problems, target search [2011_Benichou_search], and diffusion limited reactions [rice1985diffusion]. The asymptotics (6) and the relation $P_{t}(k)=P(-Q_{t}>k)$ in the semi-infinite SSEP gives a stretched exponential decay $P_{t}(k)\sim\exp{(-\sqrt{t/\tau_{s}})}$ for $k\lesssim 2\sqrt{t/\pi}$ at long-times $t$, with $\tau_{s}=[\phi(-k/\sqrt{t})]^{-2}$.
@@
The final application we discuss is about similar stretched exponential decays in kinetically constrained models [1983_Skinner_Kinetic, 1989_Spohn_Stretched, 2020_Gupta_Size, 2024_Mukherjee_Stretched], such as the energy-conserving spin-flip dynamics [1989_Spohn_Stretched, 2024_Mukherjee_Stretched], where spin auto-correlation decays as $\exp(-\sqrt{t/\tau})$, with $\tau$ remained undetermined. For the spin-model, rigorous result about $\tau$ comes [SM] from realizing that the domain-wall dynamics is equivalent to the infinite-line SSEP at a uniform density $\rho$. This correspondence relates [SM] the spin-auto-correlation to the generating function of current in the infinite SSEP leading to $\tau=[\mu_{\text{inf}}(\mathrm{i}\pi,\rho,\rho)]^{-2}$. Similar correspondence holds in presence of a slow bond or in semi-infinite geometry.
Conclusions: We derived full counting statistics (3) for SSEP on a semi-infinite line, which represents an intermediate non-equilibrium regime between the well-studied finite [2004_Bodineau_Current, 2005_Bertini_Current] and infinite SSEP [2009_Derrida_Current, 2022_Mallick_Exact]. Compared to the latter examples, the scgf (3) exhibits a piecewise solution (3b). Our exact result not only solves the absorbing boundary problem relevant to target survival [2011_Benichou_search], but also leads to the full counting statistics (24) for the infinite-SSEP with a defect bond. The variational construction used for the latter solution offers a powerful approach for studying inhomogeneous transport, even for asymmetric dynamics.
@@
There remain several related open problems of immediate interest. Most prominent among them is the scgf for fixed (quenched) initial states. The only available non-trivial result [2009_Derrida_Current2] is for the infinite line SSEP at half-filling, obtained following a relation with the fluctuating (annealed) initial state. A similar mapping for the semi-infinite SSEP yields the quenched scgf
@@
An extension of these results in higher dimensions would be interesting, where only limited results are available [2013_Akkermans_Universal, Becker_2015, Hutardo_2016, Hutardo_2017, Harris_2016]. From a practical point of view, the emergence of the SSEP in quantum circuits [2023_McCulloch_Full] or hydrodynamics in chaotic quantum systems [2024_Wienand_Emergence] offers exciting applications, where semi-infinite analog [turkeshi2024] becomes relevant. Along similar lines, extensions of our results for quantum analogues of SSEP [2019_Bernard_Open, 2021_Bernard_Can] or for integrable models [2020_Doyon_Lectures, 2022_Bastianello_Introduction] would be timely.
@@ Acknowledgements.
Acknowledgments: We acknowledge the financial support of the Department of Atomic Energy, Government of India, under Project Identification No. RTI 4002. TS thanks Bernard Derrida for insightful discussions on the problem. The importance of semi-infinite line SSEP and the additivity ansatz for slow bond originated from those discussions. TS also thanks Mustansir Barma for introducing him to the works on stretched exponential decays, particularly the open problem in [1989_Spohn_Stretched]. TS thanks the support from the International Research Project (IRP) titled ‘Classical and quantum dynamics in out of equilibrium systems’ by CNRS, France.
I End Matter
The piecewise function: The scgf (3) is plotted in Fig. 3 for parameter regimes where the piecewise nature of the function (3b) is relevant. The function $R(\omega)$ in (3c) is singular at $\omega=-1/2$, and this leads to a non-convex $R(\omega(\lambda,\rho_{a},\rho_{b}))$ as a function of $\lambda$, indicated by the dashed line in Fig. 3. The piecewise function (3b) constructed from $R(\omega)$ gives an analytic, convex scgf, which agrees with the data from rare event simulation using a cloning algorithm.
Figure 3: Scgf for semi-infinite SSEP: The scgf of current for the semi-infinite SSEP for densities $(\rho_{a},\rho_{b})\equiv(0.9,0.1)$. The solid red line represents the theoretical result (3), while the data points are obtained by the cloning algorithm with $T=500$ and $N_{c}=50,000$. The dashed line represents $R(\omega(\lambda,\rho_{a},\rho_{b}))$ in (3c) as a function of $\lambda$ which differs from the solid line for the parameter regime $\omega(\lambda,\rho_{a},\rho_{b})<-1/2$, which corresponds to $\lambda\lesssim-1.0$ in the figure.
Cumulants: The scgf (3) encapsulates all cumulants of $Q_{T}$ in the large $T$ limit for the semi-infinite SSEP. While the first three cumulants were initially reported in [2023_Saha_Current], we present here the fourth cumulant for $\rho_{a}=\rho_{b}=\rho$
Infinite line SSEP with a slow bond: The scgf of current for an infinite line SSEP with a single slow bond can be obtained by treating the entire system as composed of three subsystems: two semi-infinite systems with unit hopping rates coupled to each other by a single slow bond with hopping rate $\Gamma/\sqrt{T}$. This leads to a variational formula [SM]
$\mu_{\text{inf}}^{\text{slow}}(\lambda,\rho_{a},\rho_{b})=$ $\max_{\rho_{0},\rho_{1}}\min_{\lambda_{0},\lambda_{1}}\Big[R_{\text{si}}\big(\omega(\lambda_{0},\rho_{a},\rho_{0})\big)$
$+\Gamma\,\omega(\lambda_{1}-\lambda_{0},$ $\rho_{0},\rho_{1})+R_{\text{si}}\big(\omega(\lambda-\lambda_{1},\rho_{1},\rho_{b})\big)\Big],$ (27)
where $\rho_{0(1)}$ is the density of the common site between the left (right) semi-infinite lattice and the slow bond. The analysis is similar to the discussion in [2021_Derrida_Large] for a related problem.
Figure 4: Scgf for infinite SSEP with a slow bond: The solid blue line denotes the scgf given in (24) for $\rho_{a(b)}=0.5$ and $\Gamma=1$. The blue markers denote the corresponding numerical results obtained by a continuous-time cloning algorithm with $T=500$ and $N_{c}=10^{4}$, for $\Gamma=1$. The solid red line denotes the $\Gamma\to\infty$ limit result [2009_Derrida_Current].
Figure 5: Scgf for semi-infinite SSEP with slow boundary coupling: The solid lines denote the scgf (23) for $\rho_{a(b)}=0.5$ and different values of $\Gamma$ (green for $\Gamma=1$, blue for $\Gamma=2$, and red for $\Gamma\to\infty$). The markers denote the corresponding numerical results obtained by the cloning algorithm with $T=500$ and $N_{c}=10^{4}$.
The expression (I) further simplifies using Eq. (29) of [2021_Derrida_Large] leading to the expression in (24). We confirm this theoretical result by comparing with numerical data from our rare event simulation as shown in Fig. 4.
A similar construction leads to the scgf (23) for semi-infinite SSEP with slow coupling to reservoir. A comparison of the result with numerical data is shown in Fig. 5.
A connection between SSEP and a kinetically constrained model: The energy-conserving spin-flip dynamics on a one-dimensional infinite lattice, defined in [1989_Spohn_Stretched], where each site $i$ at a given time $t$ is assigned with an Ising spin $S_{i}(t)=\pm 1$. The dynamics is such that a spin can flip only if its two nearest neighbor’s spins have opposite signs, therefore conserving the total energy (Fig. 6). This constraint ensures the conservation of number of domain walls, whose dynamics precisely mirror the particle movement in the SSEP on an infinite lattice.
This correspondence relates the spin-auto-correlation to the particle current in SSEP. We note that whenever a domain wall passes through a site $i$ (see Fig. 6), the associated Ising spin, $S_{i}$, changes its sign. If $Q_{t}$ is the net rightward flow of domain walls through the site $i$ in time $t$, then the quantity $S_{i}(0)S_{i}(t)$ takes values +1+1 or 1-1 for $Q_{t}$ being even or odd. This implies the auto-correlation
$\big<S_{i}(0)\,S_{i}(t)\big>=P_{t}(\text{even }Q_{t})-P_{t}(\text{odd }Q_{t}),$ (28)
with the probability $P_{t}(\text{even }Q_{t})$ of $Q_{t}$ being even and $P_{t}(\text{odd }Q_{t})$ for odd, respectively.
Figure 6: Dynamics of spins [1989_Spohn_Stretched] following a kinetic constraint of fixed energy: a spin can flip with unit rate only if its nearest neighbor’s spins are of opposite sign. For example, the fifth spin from left has left neighbor down and right neighbor up, so it is allowed to flip, making the domain wall move leftward by one lattice unit. The movement of domain walls (indicated by green lines) is analogous to the dynamics of particles in a SSEP.
The difference of the two probabilities in (28) relates to the generating function of current $\big<\mathrm{e}^{\lambda Q_{t}}\big$> in SSEP. Noting that $Q_{t}$ takes only integer values, we see that for $\lambda=\mathrm{i}\pi$,
$\big<\mathrm{e}^{\mathrm{i}\pi Q_{t}}\big>=\sum_{Q_{t}}P(Q_{t})\,\mathrm{e}^{\mathrm{i}\pi Q_{t}}=P_{t}(\text{even }Q_{t})-P_{t}(\text{odd }Q_{t}).$ (29)
From (28) and (29), it is evident that at large times, the spin-auto-correlation follows the asymptotics of the current statistics in SSEP. Using the results for the latter from [2009_Derrida_Current] we see that for large times $t$,
$\big<S_{i}(0)\,S_{i}(t)\big>\sim\mathrm{e}^{\sqrt{t}\mu_{\text{inf}}(\mathrm{i}\pi,\rho,\rho)}=\mathrm{e}^{-\sqrt{{t}/{\tau}}}$ (30)
with $\tau=[\mu_{\text{inf}}(\mathrm{i}\pi,\rho,\rho)]^{-2}$. Here, the density $\rho$ is the average uniform density of domain walls in the spin model [1989_Spohn_Stretched]. The exact result (30) is consistent with the bounds derived in [1989_Spohn_Stretched].